# Data Privacy Vocabulary (DPV) — Version 2.0

### v2

@@ 1 Introduction
The modern technological landscape consists of ubiquitous digital devices and services which generate vast amounts of data, which includes sensitive information that raises privacy concerns, as well as requires the protection of data from misuse and cybersecurity threats. Regulations across the globe have been developed or updated to meet this challenge, most notably the European Union’s (EU) General Data Protection Regulation (GDPR) [GDPR] in 2016, which requires specific activities to be carried out based on defined norms and requirements, and require documenting governance processes for compliance.
@@
Based on this background, the SPECIAL H2020 project established the W3C Data Privacy Vocabularies and Controls Community Group [^1] https://www.w3.org/groups/cg/dpvcg/ (DPVCG) in 2018, which developed the Data Privacy Vocabulary [^2] https://w3id.org/dpv (DPV) [panetto_creating_2019] as a machine-readable interoperable vocabulary for the exchange of legally relevant ‘metadata’. Since then, DPV has continued being a state of the art resource that is iteratively updated to match the evolving landscape of regulations and compliance requirements. It has seen high adoption in academic, industrial, and mixed settings, and has been referenced in standards.
The DPV 1.0 focused on providing a vocabulary for describing personal data activities based on the GDPR. Since then, the world has evolved at a rapid pace with new innovations and regulations – in particular those related to ‘increasing sharing of data’ such as the EU’s Data Governance Act (DGA) [DGA], as well as for ‘regulating AI’, such as the EU’s Artificial Intelligence Act (AI Act) [AIAct]. To reflect these developments as well as others in domains such as cybersecurity, the DPVCG has updated the DPV and made it capable of representing a wide variety of legally relevant information for ‘AI and digital regulations’. In this article, we present the design and development of DPV – version 2.
@@ 2.1 Information and Knowledge Modelling
DPV 1.0 was concerned with modelling the purposes, processing operations, legal bases, personal data categories, technical and organisational measures, and the roles of entities based on the GDPR, and to fill specific gaps [panetto_creating_2019]: (1) there were no ontologies to describe and exchange information about activities involving personal data; (2) there were no agreed upon vocabularies or taxonomies for representing practical applications and uses of personal data, e.g., different purposes such as Service Provision, or Legal Bases such as Consent; and (3) there were no vocabularies that align the terminology and requirements of regulations for developing and using machine-readable information.
@@ 3 Comparison with Related Work
The Open Digital Rights Language (ODRL) [ianella2007open] is a W3C standard for modelling policies and agreements which can dictate the permissions, prohibitions, and duties associated with use of data or resources. The ODRL vocabulary also provides specific concepts such as ‘obtain consent’ to support commonly used agreements. While there is some overlap between DPV and ODRL, they are complimentary in their objectives and uses. ODRL focuses solely on providing a language for representation of policies and does not focus on specific legal requirements or jurisdictions. DPV’s concepts can thus provide the necessary legal and jurisdictional relevant information within ODRL policies. This has been explored and demonstrated by existing work [esteves_odrl_2021]. The DPVCG intends to establish close collaboration with the ODRL CG through its common members by aligning concepts between DPV and ODRL, and providing guidance for using DPV as an ODRL profile. DPV has also been mentioned by the International Data Spaces Association (IDSA) as a vocabulary of interest to support the implementation of GDPR and regulations in their policies based on ODRL [bader2020international].
LegalRuleML [^3] https://docs.oasis-open.org/legalruleml/legalruleml-core-spec/1.0.0/os/legalruleml-core-spec-1.0.0-os.html is an OASIS standard providing a ‘rule interchange language’ that enables modelling and reasoning tasks based on legal arguments. Similar to ODRL, it focuses on providing a language for expression of ‘rules’ and does not provide any taxonomies or model regulations. Similar to ODRL, DPV is also complimentary to LegalRuleML as a vocabulary that can be used to support representation of legal information. There is existing work that has explored the use of LegalRuleML for modelling requirements from the text of GDPR [robaldo2020dapreco], though it has not been maintained nor been extended to support practical requirements related to implementing GDPR.
@@ 4.1 Analysis of Citations
In this Section, the adoption of DPVCG’s outputs is evaluated through a citation analysis performed over the DPV 1.0 publication [panetto_creating_2019]. In this context, 81 publications that cite DPV were found through the Google Scholar service. The gathered results underwent a review and were included in this analysis if deemed pertinent. Duplicated publications, publications without an open-access version and in languages other than English were excluded from this analysis, resulting in 76 publications to be reviewed. Table 1 presents the results of the performed publication evaluation. The publications were evaluated in relation to their use of DPV: publications that reference DPV as a state of the art resource are signalled in the Mention column, that use DPV towards an application or use case in the Use column, and that extend DPV in the Ext. column. Publications that contributed their extensions back to DPV (Contrib. column) are also marked in Table 1, as well as if the work is applied to a certain domain or sector (Domain column). The Effort column denotes the amount of work (speculated) to update the implementation to DPV 2.0, where ++ denotes more work as implementations use a pre-1.0 version and will need to update the IRIs, + denotes minor efforts to check changed concepts in DPV 2.0, and - denotes no changes. DPV 2.0 is largely compatible with DPV 1.0, which means most implementations using DPV 1.0 can update with minimal changes.
Table 1: Citation analysis of academic publications that reference DPV 1.0 [panetto_creating_2019].
@@
[calvaresi_agentbased_2020] 2020 X Health N/A
[lieber_ecodalo_2020] 2020 X Media N/A
[bonatti_machine_2020, matulevicius_method_2020, thalhath_metaprofiles_2020] 2020 X N/A
[calvaresi_personal_2020, krasnashchok_towards_2020, leone_role_2020, ryan_design_2020] 2020 X ++
[calbimonte_decentralized_2020] 2020 X Health ++
[debruyne_justtime_2020, pandit_representing_2020, ryan_common_2020] 2020 X X ++
[esteves_challenges_2021, mcdonald_evaluation_2021, flesch_investigating_2021] 2021 X Health N/A
[grunewald_tira_2021, bonatti_representing_2021, sion_overview_2021, pandit_chapter_2021, hamed_enhancing_2021] 2021 X N/A
[ryan_gdpr_2021, esteves_odrl_2021, ekaputra_semanticenabled_2021, ryan_building_2021, leone_legal_2021] 2021 X ++
[garcia_towards_2021] 2021 X Smart products +
[hickey_gdpr_2021] 2021 X X +
[human_data_2022, jesus_consent_2022, human_advanced_2022, sangaroonsilp_mining_2022, pandit_proposals_2022, rasmusen_increasing_2022, esteves_semantifying_2022] 2022 X N/A
[esteves_analysis_2022] 2022 X X N/A
[ryan_support_2022, ryan_dpcat_2022, pandit_semantic_2022, esteves_fostering_2022] 2022 X X +
[becker_secondary_2022] 2022 X Health N/A
[esteves_using_2022, debackere_policyoriented_2022, debackere_enforcing_2022, gambarelli_privaframe_2022, kurteva_making_2022] 2022 X +
[becher_contra_2022] 2022 X IoT +
[hernandez_tikd_2022] 2022 X Health +
[pandit_making_2023, bushati_what_2023, breteler_flint_2023, kurteva_relevant_2023, asgarinia_who_2023] 2023 X N/A
[gambarelli_is_2023, grunewald_enabling_2023, bailly_prototyping_2023, tang_helping_2023, esteves_towards_2023, esteves_using_2023, taheri_compliance_2023] 2023 X +
[sun_citizencentric_2023, navarrogallinad_evaluating_2023, florea_is_2023, gallinad_usable_2023] 2023 X Health +
[kurteva_smashhitcore_2023] 2023 X Smart cities +
[gupta_oppo_2023, zichichi_decentralized_2023] 2023 X Media +
[esteves_semantics_2023, pandit_towards_2023] 2023 X X +
[raza_semantic_2023] 2023 X Health +
[pandit_enhancing_2024] 2024 X Health -
[herwanto_leveraging_2024] 2024 X +
[golpayeganiAIROOntologyRepresenting2022, golpayeganiBeHighRiskNot2023, golpayeganiAICardsApplied2024] 2024 X X WIP EU AI Act ++
[panditImplementingISOIEC2024] 2024 X -
@@
In this context, DPV’s specifications were compared against other state of the art vocabularies in the data protection domain regarding their ability to represent information related to GDPR rights and duties [esteves_analysis_2022], their machine-readability, maintenance, accessibility, GDPR support and existence of compliance tools [asgarinia_who_2023], and their capacity to aid with data interoperability and adhere to the FAIR principles [thalhath_metaprofiles_2020]. In all mentioned surveys, DPV obtained a higher score compared with other existing solutions. When it comes to extensions performed over DPV, most were contributed back to DPV to be integrated into DPVCG’s outputs. Concerning work on GDPR requirements, there were proposed extensions focusing on consent [pandit_representing_2020, debruyne_justtime_2020], in particular related to the processing of electronic health record data [raza_semantic_2023], as well as on building semantic models to represent records of processing activities [ryan_common_2020, ryan_support_2022, ryan_dpcat_2022, ryan_building_2021], data protection impact assessments [pandit_semantic_2022], data breaches’ reports [pandit_towards_2023], and international data transfer notices [hickey_gdpr_2021]. Moreover, extensions focusing on GDPR’s data subject rights and exemptions to these rights [esteves_fostering_2022] and on DGA requirements [esteves_semantics_2023, esteves_towards_2023] were also contributed back to DPVCG’s outputs.
@@
In terms of applications to specific use cases, there is a body of work focusing on providing tools for auditing and GDPR compliance evaluation [ryan_design_2020, ryan_gdpr_2021, taheri_compliance_2023], on data minimisation [garcia_towards_2021], as well as on the documentation and annotation of privacy policies [leone_role_2020, leone_legal_2021, krasnashchok_towards_2020, gupta_oppo_2023] and privacy preferences [becher_contra_2022]. Moreover, ML models were trained with DPV’s taxonomies to identify personal data processing activities in code repositories [tang_helping_2023, herwanto_leveraging_2024] and textual datasets [gambarelli_privaframe_2022, gambarelli_is_2023]. DPV’s outputs were also used to model access and usage control policies [calvaresi_personal_2020, calbimonte_decentralized_2020, ekaputra_semanticenabled_2021, zichichi_decentralized_2023], and in particular applied to Solid [esteves_odrl_2021, esteves_using_2022, debackere_policyoriented_2022, debackere_enforcing_2022, florea_is_2023, bailly_prototyping_2023, esteves_using_2023] and health data-sharing use cases [sun_citizencentric_2023, pandit_enhancing_2024], as well as to describe consent records and contracts for sensor data [kurteva_smashhitcore_2023, kurteva_making_2022]. In the context of data spaces, DPV was used to provide descriptions of health data handling activities [hernandez_tikd_2022] and to create user-centric privacy interfaces [grunewald_enabling_2023, navarrogallinad_evaluating_2023, gallinad_usable_2023].
@@ 4.3 Use of DPV in Standards
ISO/IEC TS 27560:2023 is a Technical Specification (TS) describing a ‘consent record information structure’, which defines the specific information to be maintained in consent records. DPV’s consent modelling played a significant part in the development of this standard based on sharing the knowledge [^9] Harshvardhan J. Pandit was an active contributor to the development of ISO/IEC TS 27560:2023 through the National Standards Authority of Ireland (NSAI) of legal and semantic requirements regarding consent records and receipts. The annex of 27560:2023 provides examples of consent records and receipts using JSON-LD where DPV is explicitly referenced in the document and its concepts are used as to define the schema (e.g., hasPurpose and instances (e.g., Service Provision).
@@
5.1 Overview
The motivation of the Data Privacy Vocabulary (DPV) is to provide a ‘data model’ or a ‘taxonomy’ of concepts that act as a vocabulary for the interoperable representation and exchange of information about personal data and its processing. For this, the DPV specification represents an abstract model of concepts and relationships that can be implemented and applied using technologies appropriate to the use-case’s requirements.
Figure 1: Overview of concepts in DPV and their extensions
The core concepts in DPV, as shown in Figure 1, are broadly as follows:
@@ 5.1 Overview
  • Context: concepts other than the above, e.g., Necessity, Duration, Frequency.
@@
To assist newcomers in understanding the structure of DPV and how its concepts are organised — a Primer [^10] https://w3id.org/dpv/primer document has been developed. In addition, the documentation is continually refined to provide illustrative guidance and examples [^11] https://w3id.org/dpv/2.0/examples, and a searchable index [^12] https://w3id.org/dpv/2.0/search of concepts is also provided. The below example shows the use of DPV in implementing consent records as used in ISO/IEC TS 27560:2023 Privacy technologies — Consent record information structure [panditImplementingISOIEC2024], where this work was recently also presented to the EU Commission: {minted}[fontsize=]turtle @prefix dct: <http://purl.org/dc/terms/> . @prefix dpv: <https://w3id.org/dpv#> . @prefix loc: <https://w3id.org/dpv/loc#> . @prefix eu-gdpr: <https://w3id.org/dpv/legal/eu/gdpr#> . @prefix : <https://example.com/> .
:63ded36f-4acd-4f3c-991e-6cb636698523 a dpv:ConsentRecord ; dct:hasVersion "ISO-27560" ; dpv:hasIdentifier "63ded36f-4acd-4f3c-991e-6cb636698523" ; dpv:hasDataSubject "96121fde-199f-4848-8942-4436e270513a" ; dpv:hasNotice <https://example.com/notice> ; dpv:hasProcess [ a dpv:Process ; dct:title "Send Newsletters with Seasonal Offers"@en ; dpv:hasPurpose dpv:Marketing ; dpv:hasLegalBasis dpv:Consent, eu-gdpr:A6-1-a ; dpv:hasPersonalData pd:Email ; dpv:hasDataController ex:Acme ; dpv:hasProcessing dpv:Collect, dpv:Store ; dpv:hasStorageCondition [ dpv:hasLocation loc:IE ; dpv:hasDuration "P1Y"8sd:duration ; ] ; dpv:hasJurisdiction loc:EU ; dpv:hasRecipient :Beta, :Epsilon ; ] ; dpv:hasConsentStatus dpv:ConsentGiven ; dct:hasPart [ a dpv:ConsentGiven, dpv:ExplicitlyExpressedConsent ; dpv:isIndicatedAtTime "2021-05-28T12:24:00"8sd:dateTime ; dpv:hasDuration "P1Y"8sd:duration ; dpv:hasEntity "96121fde-199f-4848-8942-4436e270513a" ] .
@@
The dpv-legal extension in DPV 1.0 was a draft providing legal concepts (laws, authorities) and locations based on ISO 3166-1 codes [^14] https://www.iso.org/iso-3166-country-codes.html. In DPV 2.0 it has been split into legal and loc (locations) extensions for separation of concerns. Both legal and location extensions are provided as completed in DPV 2.0. Further, the location concepts have been aligned with EU Vocabularies [^15] https://op.europa.eu/en/web/eu-vocabularies/concept/-/resource?uri=http://publications.europa.eu/resource/authority/country/ as an example of connecting DPV locations to external vocabularies.
@@
Changes in DPV Concepts: Of the 911 concepts in DPV 2.0, 311 are new additions and 56 concepts were removed as compared to DPV 1.0. The removed concepts represent refinements [^17] E.g. the ‘Automation’ and ‘Human Involvement’ concepts underwent a major revision based on recent developments and the publication of ISO/IEC 22989:2022 AI concepts and terminology standard https://www.iso.org/standard/74296.html and moving concepts to an extension, e.g., harms and other impacts were moved to the RISK extension.

### v3

@@ 1 Introduction
The modern technological landscape consists of ubiquitous digital devices and services which generate vast amounts of data, which includes sensitive information that raises privacy concerns, as well as requires the protection of data from misuse and cybersecurity threats. Regulations across the globe have been developed or updated to meet this challenge, most notably the European Union’s (EU) General Data Protection Regulation (GDPR) [26] in 2016, which requires specific activities to be carried out based on defined norms and requirements, and require documenting governance processes for compliance.
@@
Based on this background, the SPECIAL H2020 project established the W3C Data Privacy Vocabularies and Controls Community Group [^1] https://www.w3.org/groups/cg/dpvcg/ (DPVCG) in 2018, which developed the Data Privacy Vocabulary [^2] https://w3id.org/dpv (DPV) [67] as a machine-readable interoperable vocabulary for the exchange of legally relevant ‘metadata’. Since then, DPV has continued being a state of the art resource that is iteratively updated to match the evolving landscape of regulations and compliance requirements. It has seen high adoption in academic, industrial, and mixed settings, and has been referenced in standards.
The DPV 1.0 focused on providing a vocabulary for describing personal data activities based on the GDPR. Since then, the world has evolved at a rapid pace with new innovations and regulations – in particular those related to ‘increasing sharing of data’ such as the EU’s Data Governance Act (DGA) [28], as well as for ‘regulating AI’, such as the EU’s Artificial Intelligence Act (AI Act) [27]. To reflect these developments as well as others in domains such as cybersecurity, the DPVCG has updated the DPV and made it capable of representing a wide variety of legally relevant information for ‘AI and digital regulations’. In this article, we present the design and development of DPV – version 2.
@@ 2.1 Information and Knowledge Modelling
DPV 1.0 was concerned with modelling the purposes, processing operations, legal bases, personal data categories, technical and organisational measures, and the roles of entities based on the GDPR, and to fill specific gaps [67]: (1) there were no ontologies to describe and exchange information about activities involving personal data; (2) there were no agreed upon vocabularies or taxonomies for representing practical applications and uses of personal data, e.g., different purposes such as Service Provision, or Legal Bases such as Consent; and (3) there were no vocabularies that align the terminology and requirements of regulations for developing and using machine-readable information.
@@ 3 Comparison with Related Work
The Open Digital Rights Language (ODRL) [48] is a W3C standard for modelling policies and agreements which can dictate the permissions, prohibitions, and duties associated with use of data or resources. The ODRL vocabulary also provides specific concepts such as ‘obtain consent’ to support commonly used agreements. While there is some overlap between DPV and ODRL, they are complimentary in their objectives and uses. ODRL focuses solely on providing a language for representation of policies and does not focus on specific legal requirements or jurisdictions. DPV’s concepts can thus provide the necessary legal and jurisdictional relevant information within ODRL policies. This has been explored and demonstrated by existing work [21]. The DPVCG intends to establish close collaboration with the ODRL CG through its common members by aligning concepts between DPV and ODRL, and providing guidance for using DPV as an ODRL profile. DPV has also been mentioned by the International Data Spaces Association (IDSA) as a vocabulary of interest to support the implementation of GDPR and regulations in their policies based on ODRL [2].
LegalRuleML [^3] https://docs.oasis-open.org/legalruleml/legalruleml-core-spec/1.0.0/os/legalruleml-core-spec-1.0.0-os.html is an OASIS standard providing a ‘rule interchange language’ that enables modelling and reasoning tasks based on legal arguments. Similar to ODRL, it focuses on providing a language for expression of ‘rules’ and does not provide any taxonomies or model regulations. Similar to ODRL, DPV is also complimentary to LegalRuleML as a vocabulary that can be used to support representation of legal information. There is existing work that has explored the use of LegalRuleML for modelling requirements from the text of GDPR [72], though it has not been maintained nor been extended to support practical requirements related to implementing GDPR.
@@ 4.1 Analysis of Citations
In this Section, the adoption of DPVCG’s outputs is evaluated through a citation analysis performed over the DPV 1.0 publication [67]. In this context, 81 publications that cite DPV were found through the Google Scholar service. The gathered results underwent a review and were included in this analysis if deemed pertinent. Duplicated publications, publications without an open-access version and in languages other than English were excluded from this analysis, resulting in 76 publications to be reviewed. Table 1 presents the results of the performed publication evaluation. The publications were evaluated in relation to their use of DPV: publications that reference DPV as a state of the art resource are signalled in the Mention column, that use DPV towards an application or use case in the Use column, and that extend DPV in the Ext. column. Publications that contributed their extensions back to DPV (Contrib. column) are also marked in Table 1, as well as if the work is applied to a certain domain or sector (Domain column). The Effort column denotes the amount of work (speculated) to update the implementation to DPV 2.0, where ++ denotes more work as implementations use a pre-1.0 version and will need to update the IRIs, + denotes minor efforts to check changed concepts in DPV 2.0, and - denotes no changes. DPV 2.0 is largely compatible with DPV 1.0, which means most implementations using DPV 1.0 can update with minimal changes.
Table 1: Citation analysis of academic publications that reference DPV 1.0 [67].
@@
[11] 2020 X Health N/A
[56] 2020 X Media N/A
[6, 57, 85] 2020 X N/A
[12, 50, 54, 75] 2020 X ++
[10] 2020 X Health ++
[15, 61, 78] 2020 X X ++
[17, 58, 29] 2021 X Health N/A
[40, 7, 80, 66, 42] 2021 X N/A
[76, 21, 16, 77, 55] 2021 X ++
[34] 2021 X Smart products +
[45] 2021 X X +
[47, 49, 46, 79, 62, 70, 23] 2022 X N/A
[22] 2022 X X N/A
[73, 74, 60, 19] 2022 X X +
[5] 2022 X Health N/A
[25, 14, 13, 32, 53] 2022 X +
[4] 2022 X IoT +
[43] 2022 X Health +
[63, 9, 8, 52, 1] 2023 X N/A
[33, 39, 3, 84, 18, 20, 83] 2023 X +
[82, 59, 30, 31] 2023 X Health +
[51] 2023 X Smart cities +
[41, 86] 2023 X Media +
[24, 68] 2023 X X +
[71] 2023 X Health +
[64] 2024 X Health -
[44] 2024 X +
[37, 38, 36] 2024 X X WIP EU AI Act ++
[65] 2024 X -
@@
In this context, DPV’s specifications were compared against other state of the art vocabularies in the data protection domain regarding their ability to represent information related to GDPR rights and duties [22], their machine-readability, maintenance, accessibility, GDPR support and existence of compliance tools [1], and their capacity to aid with data interoperability and adhere to the FAIR principles [85]. In all mentioned surveys, DPV obtained a higher score compared with other existing solutions. When it comes to extensions performed over DPV, most were contributed back to DPV to be integrated into DPVCG’s outputs. Concerning work on GDPR requirements, there were proposed extensions focusing on consent [61, 15], in particular related to the processing of electronic health record data [71], as well as on building semantic models to represent records of processing activities [78, 73, 74, 77], data protection impact assessments [60], data breaches’ reports [68], and international data transfer notices [45]. Moreover, extensions focusing on GDPR’s data subject rights and exemptions to these rights [19] and on DGA requirements [24, 18] were also contributed back to DPVCG’s outputs.
@@
In terms of applications to specific use cases, there is a body of work focusing on providing tools for auditing and GDPR compliance evaluation [75, 76, 83], on data minimisation [34], as well as on the documentation and annotation of privacy policies [54, 55, 50, 41] and privacy preferences [4]. Moreover, ML models were trained with DPV’s taxonomies to identify personal data processing activities in code repositories [84, 44] and textual datasets [32, 33]. DPV’s outputs were also used to model access and usage control policies [12, 10, 16, 86], and in particular applied to Solid [21, 25, 14, 13, 30, 3, 20] and health data-sharing use cases [82, 64], as well as to describe consent records and contracts for sensor data [51, 53]. In the context of data spaces, DPV was used to provide descriptions of health data handling activities [43] and to create user-centric privacy interfaces [39, 59, 31].
@@ 4.3 Use of DPV in Standards
ISO/IEC TS 27560:2023 is a Technical Specification (TS) describing a ‘consent record information structure’, which defines the specific information to be maintained in consent records. DPV’s consent modelling played a significant part in the development of this standard based on sharing the knowledge of legal and semantic requirements regarding consent records and receipts. The annex of 27560:2023 provides examples of consent records and receipts using JSON-LD where DPV is explicitly referenced in the document and its concepts are used as to define the schema (e.g., hasPurpose and instances (e.g., Service Provision).
@@
The motivation of the Data Privacy Vocabulary (DPV) is to provide a ‘data model’ or a ‘taxonomy’ of concepts that act as a vocabulary for the interoperable representation and exchange of information about personal data and its processing. For this, the DPV specification represents an abstract model of concepts and relationships that can be implemented and applied using technologies appropriate to the use-case’s requirements. The core concepts in DPV, as shown in Figure 1, are broadly as follows:
@@
  • Context: concepts other than the above, e.g., necessity, duration
@@
To assist newcomers in understanding the structure of DPV and how its concepts are organised — a Primer [^9] https://w3id.org/dpv/primer document has been developed. In addition, the documentation is continually refined to provide illustrative guidance and examples [^10] https://w3id.org/dpv/2.0/examples, and a searchable index [^11] https://w3id.org/dpv/2.0/search of concepts is also provided. The below example shows the use of DPV in implementing consent records as used in ISO/IEC TS 27560:2023 Privacy technologies — Consent record information structure [65], where this work was recently also presented to the EU Commission: {minted}[fontsize=]turtle @prefix dct: <http://purl.org/dc/terms/> . @prefix dpv: <https://w3id.org/dpv#> . @prefix loc: <https://w3id.org/dpv/loc#> . @prefix eu-gdpr: <https://w3id.org/dpv/legal/eu/gdpr#> . @prefix : <https://example.com/> . :63ded36f-4acd-4f3c-991e-6cb636698523 a dpv:ConsentRecord ; dct:hasVersion "ISO-27560" ; dpv:hasIdentifier "63ded36f-4acd-4f3c-991e-6cb636698523" ; dpv:hasDataSubject "96121fde-199f-4848-8942-4436e270513a" ; dpv:hasNotice <https://example.com/notice> ; dpv:hasProcess [ a dpv:Process ; dct:title "Send Newsletters with Seasonal Offers"@en ; dpv:hasPurpose dpv:Marketing ; dpv:hasLegalBasis dpv:Consent, eu-gdpr:A6-1-a ; dpv:hasPersonalData pd:Email ; dpv:hasDataController ex:Acme ; dpv:hasProcessing dpv:Collect, dpv:Store ; dpv:hasStorageCondition [ dpv:hasLocation loc:IE ; dpv:hasDuration "P1Y"8sd:duration ; ] ; dpv:hasJurisdiction loc:EU ; dpv:hasRecipient :Beta, :Epsilon ; ] ; dpv:hasConsentStatus dpv:ConsentGiven ; dct:hasPart [ a dpv:ConsentGiven, dpv:ExplicitlyExpressedConsent ; dpv:isIndicatedAtTime "2021-05-28T12:24:00"8sd:dateTime ; dpv:hasDuration "P1Y"8sd:duration ; dpv:hasEntity "96121fde-199f-4848-8942-4436e270513a" ] .
@@
The dpv-legal extension in DPV 1.0 was a draft providing legal concepts (laws, authorities) and locations based on ISO 3166-1 codes. In DPV 2.0 it has been split into legal and loc (locations) extensions for separation of concerns. Both legal and location extensions are provided as completed in DPV 2.0. Further, the location concepts have been aligned with EU Vocabularies [^13] https://op.europa.eu/s/zQhL as an example of connecting DPV locations to external vocabularies.
@@
Changes in DPV Concepts: Of the 911 concepts in DPV 2.0, 311 are new additions and 56 concepts were removed as compared to DPV 1.0. The removed concepts represent refinements and moving concepts to an extension, e.g., harms and other impacts were moved to the RISK extension.